The two common chlorides of phosphorus, PCl3 and PCl5, both important in the production of other phosphorus compounds, coexist in equilibrium through the reaction PCl3(g) + Cl2(g) → PCl5(g).


At 250oC, an equilibrium mixture in a 2.50 L flask contains 0.105 g PCl5, 0.220 g PCl3, ande 2.12 g Cl2. What are the values of (a) Kc and (b) Kp for this reaction at 250oC? Write the full computation of your answer.


1
Expert's answer
2021-04-28T02:46:17-0400

Solution:

The molar mass of PCl5 is 208.24 g/mol.

The molar mass of PCl3 is 137.33 g/mol.

The molar mass of Cl2 is 70.906 g/mol.


Convert grams to Molarity (mol/L):

Molarity of PCl5 = (0.105 g PCl5 / 2.50 L) × (1 mol PCl5 / 208.24 g PCl5) = 0.0002017 M

[PCl5] = 0.0002017 M

Molarity of PCl3 = (0.220 g PCl3 / 2.50 L) × (1 mol PCl3 / 137.33 g PCl3) = 0.0006408 M

[PCl3] = 0.0006408 M

Molarity of Cl2 = (2.120 g Cl2 / 2.50 L) × (1 mol Cl2 / 70.906 g Cl2) = 0.0119595 M

[Cl2] = 0.0119595 M


Write out the reaction:

PCl3(g) + Cl2(g) → PCl5(g)

Write out the expression for Kc:



Plug in the Molarity given:




Kc = 26.319 M-1


Kp = Kc × (RT)Δn

Δn = 1 - (1+1) = -1

T = 250oC + 273.15 = 523.15 K

R = 0.08206 L atm mol-1 K-1

Hence,

Kp = (26.319) × (0.08206 × 523.15)-1 = 0.61307 = 0.613

Kp = 0.613 atm-1


Answers:

(a) Kc = 26.319 M-1

(b) Kp = 0.613 atm-1